Active mass of water at 4 oC .how can we calculate the 55.5

Grade:12th pass

1 Answers

Arun
6 years ago
For water, 
density = 1 g/mL = 1000 g/L 
Molar mass = 18.015 g/mol 

so 

1000 g/L = (1000 g/L)/(18.015 g/mol) = 55.51 mol/L